# Faster and Better 3D Splatting via Group Training

### v1

@@ 3 Methodology
In Sec. 3.1, we will provide a comprehensive introduction to the Group Training method and examine its diverse functions across training 3D Gaussian Splatting. Subsequently, we introduce two different sampling strategies: Random Sampling and Prioritized Sampling (including important score, opacity-based, volume-based.) in Sec. 3.2, and we compare the computational efficiency and performance loss of various Prioritized Sampling. The impact of different Prioritized Sampling strategies on gaussian densification and rendering is also examined. Finally, in Sec. 3.3 , we summary training characteristics of 3DGS with Group Training, and the Sampling Strategies are identified.
@@ 3.1 Group Training
As previously outlined in Sec. 2, the 3DGS training process consists of two distinct phases: “Densification” and “Optimization.” Group Training is facilitated in both phases and concludes with Groups Merging at each phases, as illstrated in Fig. 4. Groups Merging without resampling is conducted at iteration $I_{\text{D}}$ and $I_{\text{O}}$. Groups Merging requires minimal time to integrate all Gaussian primitives for global scene reconstruction. The Groups Merging for global densification also aids in removing redundant Gaussians. The detailed analysis is provided in Sec. 4.3.
The effects of Group Training vary across these two phases. When group training is enabled during Gaussian densification, there is a notable reduction in the number of Gaussians within the space. This reduction significantly impacts the densification process, as will be examined in the following experiments. Further details, along with an analysis of the results, are provided in Sec. 4.2.
3.2 Efficient and Effective Sampling Strategies
In each grouping, a proportion of Gaussians is sampled from the merged Gaussians. The simplest sampling strategy is Random Sampling (RS), which selects Gaussians uniformly at random. To enhance the training process of the 3DGS and improve rendering quality, we propose a Prioritized Sampling approach that probabilistically selects Gaussians based on their attributes.
We evaluated various prioritized sampling strategies, including opacity and volume-based approaches [5] and importance scores [25, 6, 32, 8]. Among these, Opacity-based Prioritized Sampling demonstrated superior performance in both training speed and rendering quality. While importance scores effectively evaluate primitive significance, they suffer from computational overhead and stage-dependent reliability, particularly during early training. Based on this analysis, we examine the effectiveness of different sampling strategies through the lens of Gaussian densification characteristics and rendering formulation (detailed comparisons are provided in the appendix).
@@ 3.2.1 Sampling Strategy for Better Densification
Gaussian densification constitutes a pivotal element in the enhancement of the scene representation capabilities of 3DGS. The sources Gaussians lead to densification were tracked, with particular attention paid to the opacity values and volumes of Gaussian primitives that were cloned or split. Subsequently, a statistical analysis was conducted on the distribution of Gaussian attributes during each densification step, as illustrated in Fig. 5.
It is evident that Gaussians with high opacity and small volume are the primary sources implementing Gaussian densification. The Gaussians that contribute to densification are predominantly situated around 1. As densification progresses, Gaussians with lower volume become increasingly involved in the densification process. Notably, a lack of such Gaussians leads to an enhancement in under-reconstruction and over-reconstruction [12], thereby facilitating the gaussian densification. This typically leads to an increase in the number of redundant Gaussians. Nevertheless, while small volumes can be employed as a compression metric [5] in the model compression of 3D Gaussian Splatting, they also represent a primary source of density during the densification process. Therefore, volume is not a suitable metric for Prioritised Sampling. As a result, it is advisable to cache Gaussians with high opacity for better gaussian densification .
@@
Figure 5: The distribution of Gaussian attributes. The distribution of all Gaussian attributes (left) and those contributing specifically to densification (right) in the “Bicycle”[1] scene reconstructed with 3DGS [12]. Top row: The distribution of Gaussian Opacity, which is primarily concentrated around 0 and 1. In contrast, the Gaussians that contribute to densification are predominantly situated around 1. Bottom row: The distribution of Gaussian Volume. As densification progresses, Gaussians with lower volume become increasingly involved in the densification process.
@@ 3.2.2 Sampling Strategy for Faster Rendering
The proposal to speed up training by limiting the number of Gaussians is obvious. A more detailed examination revealed that the distribution of opacity has a significant impact on the computational efficiency of the entire rendering framework. To render an image from a specific viewpoint in the 3DGS model, the framework employs the $\alpha$-blending [20] formula repeatedly when rendering any given pixel, $p$.
$C_{p}=\sum_{i=1}^{N}c_{i}\alpha_{i}T_{i},\quad T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$ (3)
where $N$ represents the number of Gaussians that are necessary to render pixel $p$, which is also the number of Gaussians that are being optimised. In essence, $N$ represents the primary determinant of rendering speed; a reduction of $N$ will directly enhance rendering speed. The sole criterion that determines $N$ is the saturation of accumulated opacity $\alpha$ in a pixel in 3DGS [12]. In particular, this signifies that $T_{i}$ is below the specified threshold, which has been preset to 0.0001. It is evident that the presence of high opacity can accelerate the saturation of accumulated opacity $\alpha$, while an excess of low-opacity Gaussians significantly increases the computational burden of rendering. In other words, the distribution of opacity can influence the speed at which saturation of $\alpha$ occurs, thereby affecting rendering speed.
@@ 3.3 Analysis and Design Insights
High-opacity Gaussians have been shown to play a crucial role throughout the entire training process. During Gaussian densification, increasing the sampling probability of Under-training high-opacity Gaussians helps minimize the creation of redundant points. Additionally, high-opacity Gaussians contribute to faster saturation of accumulated opacity by constraining the opacity distribution of Under-training Gaussians.
Through a comprehensive testing program and a detailed examination of the densification and rendering formulas in 3DGS, we have identified the acceleration effect of the opacity attribute in the 3DGS training process. Based on these insights, we propose an Opacity-based Prioritised Sampling (OPS) strategy. The Opacity-based sampling probability $p_{i}$ for selecting an Under-training Gaussian $G_{i}$ is defined as:
@@ 4 Experiments
We undertake an experimental investigation of Group Training within the context of 3D Gaussian Splatting [12] and Mip-Splatting [30]. The experimental settings are detailed in Sec. 4.1. Subsequently, we present the experimental results in Sec. 4.2. The ablation experiments and their results are comprehensively discussed in Sec. 4.3. Finally, we analyse the results of Opacity-based Prioritised Sampling (OPS) and Group Training in  Sec. 4.4.
@@ 4.1 Experimental Settings
Implementation Details. The Group Training method was integrated as a plug-in to facilitate its utilisation in conjunction with 3DGS. Specifically, Group Training was incorporated before rendering each training iteration, with its activation controlled by the iteration index. We implemented our method within both 3D Gaussian Splatting [12] [^1] https://github.com/graphdeco-inria/gaussian-splatting and Mip-Splatting [30] [^2] https://github.com/autonomousvision/mip-splatting.
Group Training was implemented with a default grouping interval of 500 iterations throughout the training process. We employed both Random Sampling (RS) and Opacity-Based Prioritized Sampling (OPS) strategies with an Under-Training Ratio (UTR) of 0.6. To ensure computational stability, considering the importance of initial Gaussians [4], Group Training was activated at the 500th iteration after densification begins. The global densification ($I_{\text{D}}$) and optimization ($I_{\text{O}}$) were scheduled at 14.5K and 29K iterations, respectively.
To manage the increased GPU memory consumption during group merging, we fixed the SH coefficients [26, 24] at iterations $I_{\text{D}}$ and $I_{\text{O}}$. We also conducted experiments applying Group Training solely during the densification phase. Complete experimental results are presented in Sec. 4.2.
Dataset and Metrics. The efficacy of our method is evaluated on the Mip-NeRF360 [1] [^3] https://jonbarron.info/mipnerf360/, Tanks & Temples [13] [^4] https://www.tanksandtemples.org/, Deep Blending [9] [^5] http://visual.cs.ucl.ac.uk/pubs/deepblending/, and NeRF-Synthetic [20] [^6] https://www.matthewtancik.com/nerf datasets using a single GTX RTX 3090. Subsequently, the model’s reconstruction performance is evaluated using PSNR, SSIM, and LPIPS. Furthermore, the training time, peak memory usage, and model size for each scene are recorded.
@@ 4.2 Experiments Results
This section presents the experimental results of integrating the Group-Training method into the 3D Gaussian Splatting [12] and Mip-Splatting [30]. The test outcomes demonstrate that the Group Training with OPS offers a significant reduction in training time compared to the baseline in all reconstructed scenes, while concurrently enhancing the rendering quality of the novel view. A comprehensive visualization of the contrasts is presented in Fig. 6.
3D Gaussian Splatting. We first tested the effectiveness of Group Training based on 3DGS. The reconstruction efficiency for all tested scenes is summarized in Tab. 1, and Tab. 2 presents the novel view synthesis rendering quality across all reconstructed scenes.
Across these scenes, enabling Group Training consistently resulted in faster reconstruction, and this speedup was more pronounced in complex scenes. In addition to accelerating scene reconstruction, Group Training also steadily improved the quality of reconstruction.
However, we found that Group Training with RS had an unstable effect on model size. Randomly sampling Gaussian primitives to cache can lead to an overabundance of redundant Gaussian primitives in some reconstruction tasks, which complicates subsequent operations, such as model compression [24, 23, 26]. Additionally, GPU memory usage significantly increased due to this redundancy [16], as shown in Tab. 1. In comparison to Group Training with RS, Group Training with OPS demonstrated a more pronounced improvement in reconstruction speed. For the reconstruction of all three complex scenes, the average speed increase was about 30%, and for the reconstruction of simpler scenes, the speed increase was about 20%. Moreover, Group Training with OPS greatly reduced the generation of redundant Gaussian primitives, leading to a smaller reconstruction model size than the original 3DGS reconstruction model in all tasks. Across all reconstruction tasks, the model size was reduced by between 10% and 40%, resulting in a more compact model overall.
Mip-Splatting. We tested Group Training with Mip-Splatting, another significant variant of the Gaussian Splatting model. The experimental results for Mip-Splatting are presented in Tab. 3 and Tab. 4. Overall, the performance of Group Training with Mip-Splatting was similar to that observed with 3DGS. In our experiments, Group Training with RS consistently improved both reconstruction quality and speed for Mip-Splatting but produced the highest number of Gaussian primitives during the densification phase. Group Training with OPS achieved the fastest reconstruction speed, and its reconstruction quality was nearly optimal. Moreover, the OPS strategy reduced the proportion of densification and resulted in a smaller model size that was either the smallest among the tested methods or comparable to the original Mip-Splatting model.
@@
    Group Training with OPS simultaneously achieves better reconstruction quality with reduced model size. This demonstrates that superior performance can be achieved through more effective primitive organization rather than simply increasing model capacity.
These findings indicate that the benefits of Group Training stem from its fundamental impact on the optimization process rather than from model size manipulation. We provide a detailed analysis of these phenomena in Sec. 4.4.
@@ 4.3 Ablation Studies
The results indicate that cyclic resampling contributes the most to improving training speed. Specifically, during iterations $I_{\text{D}}$$\sim$15K, the Global Densification process adaptively performs cyclic pruning of low-opacity Gaussians within 3DGS, reducing redundant Gaussians, decreasing the model size, and thereby enhancing training efficiency. The effectiveness of Global Optimization during iterations $I_{\text{O}}$ $\sim$ 30K depends on whether Global Densification is enabled. When Global Densification is not used, the Caching Group retains a high proportion of redundant Gaussians, leading to an increase in redundancy if Global Optimization is applied under these conditions. In contrast, when Global Densification is enabled, the 3DGS model removes most redundant Gaussians during global densification, resulting in fewer redundant Gaussians in the Caching Group during the Optimization phase. In this scenario, Global Optimization can make full use of all Gaussians, further improving reconstruction performance.

### v2

@@ 3 Methodology
In Sec. 3.1, we will provide a comprehensive introduction to the Group Training method and its multifunctional advantages in 3D Gaussian Splatting (3DGS). Section 3.2 introduces two sampling paradigms: Random Sampling and Prioritized Sampling (incorporating importance score, opacity-based, and volume-based criteria), with comparative analysis of their computational efficiency and performance trade-offs. We mathematically demonstrate that Opacity-based Prioritized Sampling achieves effective 3DGS densification and efficient rendering. Finally, in Sec. 3.3 , we summary training characteristics of 3DGS with Group Training, and the Sampling Strategies are identified.
@@ 3.1 Group Training
As outlined in Sec. 2, the 3DGS training process consists of two distinct phases: “Densification” and “Optimization”. Group Training is facilitated in both phases and concludes with Groups Merging at each phases, as illstrated in Fig. 4.
3.2 Effective and Efficient Sampling Strategies
Within each groupinsg, a subset of Gaussians is sampled from the merged Gaussians. The simplest sampling strategy is Random Sampling (RS). We propose attribute-driven Prioritized Sampling to optimize 3DGS further.
We evaluated various attributes , including opacity and volume-based approaches [5] and importance scores [25, 6, 32, 8] (see appendix for detailed comparisons). Among these, Opacity-based Prioritized Sampling demonstrated superior performance in both training speed and rendering quality. As mathematically proven in  Sec. 3.2.1 and 3.2.1, opacity intrinsically governs two critical 3DGS properties, both experimentally validated.
@@ 3.2.1 Sampling Strategy for Better Densification
For any Gaussian primitive $G_{m}$ projected onto the imaging plane with 2D center coordinates $[x_{\mathrm{m}},y_{\mathrm{m}}]$. The partial derivatives of loss $L$ w.r.t the projected coordinates are $[\frac{\partial L}{\partial x_{\mathrm{m}}},\frac{\partial L}{\partial y_{\mathrm{m}}}]$, which govern the densification criterion by:
$\sqrt{\left(\frac{\partial L}{\partial x_{\mathrm{m}}}\right)^{2}+\left(\frac{\partial L}{\partial y_{\mathrm{m}}}\right)^{2}}>\tau_{\mathrm{grad}},$
where $\tau_{\mathrm{grad}}$ denotes the predefined gradient threshold.
We formulate the first fundamental property of 3DGS:
$\bullet$ Proposition 1. Opacity-based Effective Gaussians Densification: Gaussian primitives with higher opacity serve as the primary contributors to densification of 3DGS.
Proof. Under our assumptions of mutual independence between Gaussian attributes and within-primitive parameter independence, the partial derivatives are computed as:
$\frac{\partial L}{\partial x_{\mathrm{m}}}&=o_{m}\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial\hat{C}}\frac{\partial\hat{C}}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x}\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\end{aligned}$
$\frac{\partial L}{\partial y_{\mathrm{m}}}&=o_{m}\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial\hat{C}}\frac{\partial\hat{C}}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta y}\frac{\partial\Delta y}{\partial y_{\mathrm{m}}},\end{aligned}$
where $o_{m}$ denotes the opacity of $G_{m}$, $\hat{C}$ represents the rendered pixel value, and $\Delta x,\Delta y$ indicate coordinate offsets between pixels and $[x_{\mathrm{m}},y_{\mathrm{m}}]$. Under the parameter mutual independence assumptions, the $[\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\frac{\partial\Delta y}{\partial y_{\mathrm{m}}}]$, $[\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x},\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta y}]$ and $\frac{\partial L}{\partial\hat{C}}$ remain independent of $o_{m}$. We prove that $\frac{\partial\hat{C}}{\partial\alpha_{m}}$ increases with higher expected value of $o_{m}$ in appendix:
$\mathbb{E}\left[\frac{\partial\hat{C}}{\partial\alpha_{m}}\right]=\frac{(c_{0}-c_{\mathrm{bg}})T_{\mathrm{saturation}}}{1-\mathbb{E}[o_{m}]\cdot\mathbb{E}[G_{m}^{\mathrm{2D}}]},$
where $c_{0}$ and $\alpha_{0}$ denote the expected color and opacity respectively, and $T_{\mathrm{saturation}}$ controls $\alpha$ saturation [12].
With fixed $L$Eqs. 4, 5 and 6 demonstrate that elevated opacity distributions yield greater gradient magnitudes in $\left[\frac{\partial L}{\partial x_{\mathrm{m}}},\frac{\partial L}{\partial y_{\mathrm{m}}}\right]$, thereby making such Gaussians statistically prioritized for densification. $\Box$
@@
Figure 5: The distribution of Gaussian attributes. The distribution of all Gaussian attributes (left) and those contributing specifically to densification (right) in the “Bicycle”[1] with 3DGS [12]. Top row: While the opacities are primarily concentrated around 0 and 1, the Gaussians that contribute to densification are predominantly situated around 1. Bottom row: The distribution of Volume. As densification progresses, Gaussians with lower volume become increasingly involved in the densification process.
$\bullet$ Experimental Validation: Our experiments track Gaussian primitives contributing to densification, focusing on opacity and volumes values. Statistical analysis of attribute distributions during densification steps (Fig. 5) reveals that Gaussians with high opacity and small volume are the primary sources implementing Gaussian densification.
Notably, insufficient high-opacity Gaussians increase the photometric loss, exacerbating both under-reconstruction and over-reconstruction [12], thereby forces redundant densification. While small volumes can be employed as a compression metric [5] in the model compression of 3D Gaussian Splatting, they simultaneously drive reasonable densification. Take into account all factors, it is advisable to cache Gaussians with low-opacity for better gaussian densification.
@@ 3.2.2 Sampling Strategy for Faster Rendering
During viewpoint-specific image synthesis in 3DGS, the framework implements $\alpha$-blending [20] through recursive application of the compositing equation on a per-pixel.
$C_{p}=\sum_{i=1}^{N}c_{i}\alpha_{i}T_{i},\quad T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$
$N=\min\left\{i\in\mathbb{N}^{+},\bigg{|}\prod_{j=1}^{i-1}(1-\alpha_{j})\leq T_{\mathrm{saturation}}\right\},$
where the parameter $N$ denotes the number of Gaussians required to render pixel $p$. As established in 3DGS [12], the determination of $N$ hinges on achieving opacity saturation within a pixel’s accumulated $\alpha$ value (Transmittance $T_{i}<T_{\mathrm{saturation}}$). This critical threshold implementation introduces a fundamental constraint on rendering efficiency: during forward rendering operations, the sequential traversal of Gaussian primitives inherently prevents parallel computation of $T_{i}$ values. Consequently, $N$ emerges as the principal determinant of rendering speed, where any reduction on $N$ directly translates to accelerated renderings.
Based on the preceding analysis, we formulate the second fundamental property of the 3DGS framework:
$\bullet$ Proposition 2. Opacity-based Efficient Rendering Acceleration: Gaussian primitives with higher opacity enable faster rendering through faster achieving $\alpha$ saturation.
Proof. Building upon the premise of mutually independent $\alpha_{i}$ values between Gaussians, we posit that the mathematical expectation $\mathbb{E}[T_{N}]\approx T_{\mathrm{saturation}}$,
$\mathbb{E}[T_{N}]=(1-\mathbb{E}\left[{\alpha_{i}}\right])^{N}=(1-\mathbb{E}[o_{i}]\cdot\mathbb{E}[G_{i}^{\mathrm{2D}}])^{N},$ (9)
where $\alpha_{i}$ combines its intrinsic opacity $o_{i}$ and $G_{i}^{\mathrm{2D}}$ (dependent on view projection and covariance, independent of $o_{i}$). We decouple the expected value of $\alpha_{i}$. $\Box$
$\bullet$ Experimental Validation: We generated a 3DGS model with Gaussian $o_{i}$ values sampled from $\mathcal{N}(\mu_{o},0.1)$, fixing the other parameters. The rendering time as a function of $\mu_{o}$ is shown in Fig. 6. The results reveals that Gaussian primitives with higher-opacity induce faster transmittance saturation, thereby reducing the blending steps $N$ and rendering computational load.
@@ 3.3 Analysis and Design Insights
Through a comprehensive testing program and a detailed examination of the densification and rendering formulas in 3DGS, we have identified the acceleration effect of the opacity attribute in the 3DGS training process. Based on these insights, we propose an Opacity-based Prioritised Sampling (OPS) strategy.
The Opacity-based sampling probability $p_{i}$ for selecting an Under-training Gaussian $G_{i}$ is defined as:
@@
Refer to captionRefer to caption
Figure 6: High-opacity Gaussians accelerate 3DGS rendering. Left: The rendering time decreases by 40% as $\mu_{o}$ increases. Right: The mechanisms that highers-opacity Gaussians achieve faster $\alpha$ saturation with less number of Gaussians.
@@ 4 Experiments
We undertake an experimental investigation of Group Training within the context of 3D Gaussian Splatting [12] and Mip-Splatting [30]. The experimental settings are detailed in Sec. 4.1. Subsequently, we present the experimental results in Sec. 4.2. Cross-baselines applicability validation across diverse scenarios is further presented in Sec. 4.3. The ablation experiments and their results are comprehensively discussed in Sec. 4.4. Finally, we analyse the results of different sampling strategies in  Sec. 4.5.
@@ 4.1 Experimental Settings
Implementation Details. The Group Training method was integrated as a plug-in to facilitate its utilisation in conjunction with 3DGS. We implemented our method within both 3D Gaussian Splatting [12] [^1] https://github.com/graphdeco-inria/gaussian-splatting and Mip-Splatting [30] [^2] https://github.com/autonomousvision/mip-splatting.
Group Training was implemented with a default grouping interval of 500 iterations throughout the training process. We employed both Random Sampling (RS) and Opacity-Based Prioritized Sampling (OPS) strategies with an Under-Training Ratio (UTR) of 0.6. To ensure computational stability, considering the importance of initial Gaussians [4], Group Training was activated at the 500th iteration after densification begins. The global densification ($I_{\text{D}}$) and optimization ($I_{\text{O}}$) were scheduled at 14.5K and 29K iterations, respectively. To manage the increased GPU memory consumption during group merging, we fixed the SH coefficients [26, 24] at iterations $I_{\text{D}}$ and $I_{\text{O}}$. We also conducted experiments applying Group Training solely during the densification phase.
Dataset and Metrics. The efficacy of our method is evaluated on the Mip-NeRF360 [1] [^3] https://jonbarron.info/mipnerf360/, Tanks & Temples [13] [^4] https://www.tanksandtemples.org/, Deep Blending [9] [^5] http://visual.cs.ucl.ac.uk/pubs/deepblending/, and NeRF-Synthetic [20] [^6] https://www.matthewtancik.com/nerf using a single RTX 3090. Subsequently, the model’s reconstruction performance is evaluated using PSNR, SSIM, and LPIPS. Furthermore, the training time, peak memory usage, and model size for each scene are recorded.
@@ 4.2 Experiments Results
This section presents the experimental results of integrating the Group-Training method into the 3D Gaussian Splatting [12] and Mip-Splatting [30]. The test outcomes demonstrate that the Group Training with OPS offers a significant reduction in training time compared to the baseline in all scenes, while concurrently enhancing the rendering quality of the novel view, as presented in Fig. 7.
3D Gaussian Splatting. We first tested the effectiveness of Group Training based on 3DGS. The reconstruction efficiency for all scenes is summarized in Tab. 1, and Tab. 2 presents the NVS rendering quality across all scenes.
Across these scenes, Group Training consistently resulted in faster reconstruction, and this speedup was more pronounced in complex scenes. However, Group Training with RS could lead to an overabundance of redundant Gaussian primitives in some reconstruction tasks, which complicates subsequent operations, such as model compression [24, 23, 26]. Additionally, GPU memory usage significantly increased due to this redundancy [16], as shown in Tab. 1. In comparison to Group Training with RS, Group Training with OPS demonstrated a more pronounced improvement in reconstruction speed. For the reconstruction of all three complex scenes, the average speed increase was about 30%. Moreover, OPS greatly reduced the generation of redundant Gaussian primitives, leading to a smaller reconstruction model size than the original 3DGS reconstruction model in all tasks. Across all reconstruction tasks, the model size was reduced by between 10% and 40%, resulting in a more compact model overall.
Mip-Splatting. We tested Mip-Splatting with Group Training, another significant variant of the Gaussian Splatting model. The experimental results are presented in Tab. 3 and Tab. 4. Overall, the performance of Group Training with Mip-Splatting was similar to 3DGS. Group Training with RS consistently improved both reconstruction quality and speed for Mip-Splatting but produced the highest number of Gaussian primitives during the densification phase. Group Training with OPS achieved the fastest reconstruction speed, and its reconstruction quality was nearly optimal. The OPS reduced the proportion of densification and resulted in a smaller model size that was either the smallest among the tested methods or comparable to the baseline.
@@
    Group Training with OPS simultaneously achieves better reconstruction quality with reduced model size. This demonstrates that superior performance stem from its fundamental impact on the optimization process rather than from model size.
4.3 Methodological Applicability
Group-Training framework fundamentally addresses accelerated training and rendering in 3DGS through its dataset-agnostic improvement. Designed as a pluggable training component, the method demonstrates dual efficacy in both 3DGS acceleration model [6] and compression model [5], which is detailed in appendix.
@@ 4.4 Ablation Studies
The results indicate that cyclic resampling contributes the most to improving training speed. Specifically, during iterations $I_{\text{D}}$$\sim$15K, the Global Densification process adaptively performs cyclic pruning of low-opacity Gaussians within 3DGS, reducing redundant Gaussians, decreasing the model size, and thereby enhancing training efficiency. When Global Densification is enabled, the 3DGS model removes most redundant Gaussians during global densification, resulting in fewer redundant Gaussians in the Caching Group during the Optimization phase. In this scenario, Global Optimization can make full use of all Gaussians, further improving reconstruction performance.